27) In a raffle, one ticket costs $3.00. There is a 0.002 chance of winning $500, a 0.004 chance of winning $200, and a 0.006 chance of winning $100. Find the expected profit from buying one ticket? Create an appropriate table to answer this question.
Answer #1

SOLUTION-

COST OF EACH TICKET= $3.00

PROBABILITY OF WINNING $500=0.002

PROBABILITY OF WINNING $200=0.004

PROBABILITY OF WINNING $100=0.006

AS, SUM OF PROBABILITIES MUST BE 1, PROBABILITY OF WINNING $0 = 1-(0.002+0.004+0.006)

= 1-0.012 = 0.988

Expected Wins= (500*0.002)+(200*0.004)+(100*0.006)

= $2.40

Hence, Expected Profit= -3.00+2.40

= -$ 0.60

**REMARK**- IN CASE OF DOUBT, COMMENT BELOW. AND IN CASE YOU LIKED THE SOLUTION, DO LIKE